OpenStudy (anonymous):

The sum of two numbers is 88. One number is 20% more than the other. Find the numbers.

OpenStudy (anonymous):

x + 1.2 x = 88 {40,48}

OpenStudy (zzr0ck3r):

x+y = 88 x = y + .20(y) so y+(.20)y + y = 88 2y+.20y = 88 y(2+.20) = 88 y = 88/2.20 = 40 so 40 + x = 88 thus x = 48